Use the Pythagorean theorem to determine the length of each line segment. Where appropriate, round to the nearest hundredth.
48.) (3, 2), (1, -1)

Answer by checkley71(8403)

THE X DISTANCE=3-1=2
THE Y DISTANCE=2+1=3
2^2+3^2=H^2 (H IS THE HYPOTENUSE OR THE DISTANCE BETWEEN POINTS)
4+9=H^2
H^2=13
H=SQRT13
H=3.6 DISTANCE BETWEEN THESE 2 POINTS.